OpenStudy (anonymous):

Please help, I want guidance not answers please work through it with me, medal to best guide :) Jose is paying for an eleven dollar meal using bills in his wallet. He has four one dollar bills, two five dollar bills, and two ten dollar bills. If he selects two bills at random, one at a time from his wallet, what is the probability that he will choose a ten dollar bill and a one dollar bill to pay for the meal? Show your work.

jimthompson5910 (jim_thompson5910):

focus on the first part: what is the probability he will choose a 10 dollar bill (ignore the second selection for now)

jimthompson5910 (jim_thompson5910):

let me know if you can do this part at all

jimthompson5910 (jim_thompson5910):

or tell me if you're completely stuck with everything

OpenStudy (anonymous):

I'm so lost...

OpenStudy (anonymous):

He has 4 one dollar bills, two five dollar bills and two ten dollar bills.. If he picks two bills... What are all the possibilities.. \[\frac{ 8! }{ 4!2!2! }\] There is the 4! because we have a repeat of dollar bills and 2!'s for the two fives and two tens.

jimthompson5910 (jim_thompson5910):

ok can you tell me how many bills there are total?

jimthompson5910 (jim_thompson5910):

still stuck? or did you figure it out?

OpenStudy (anonymous):

Sorry I get side tracked a lot i'm trying to focus hold on

jimthompson5910 (jim_thompson5910):

you're ok, take all the time you need

OpenStudy (anonymous):

$34 is total and bills in total are 8 right?

jimthompson5910 (jim_thompson5910):

yep there are 8 bills in total

jimthompson5910 (jim_thompson5910):

of these 8 bills, how many are tens?

OpenStudy (anonymous):

2, I'm sorry dyslexia and math don't go together

jimthompson5910 (jim_thompson5910):

that's ok, keep at it and you'll get better

jimthompson5910 (jim_thompson5910):

so the probability of picking a ten is 2/8 = 1/4

OpenStudy (anonymous):

Okay. I get that.

jimthompson5910 (jim_thompson5910):

once you pick out a ten, you'll have 8 - 1 = 7 bills left over

OpenStudy (anonymous):

Otay..

jimthompson5910 (jim_thompson5910):

now how many one dollar bills are there

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

so the probability of picking out a one dollar bill is 4/7 (this is after you picked a ten first)

jimthompson5910 (jim_thompson5910):

you then multiply it with 1/5 to get (1/5)*(4/7) = (1*4)/(5*7) = 4/35

jimthompson5910 (jim_thompson5910):

so the probability of picking a ten, then picking a one is 4/35

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

no....whered the 35 come from?

jimthompson5910 (jim_thompson5910):

5*7 = 35

jimthompson5910 (jim_thompson5910):

oh wait, sorry

jimthompson5910 (jim_thompson5910):

mixed up numbers, my bad

jimthompson5910 (jim_thompson5910):

it should be (1/4)*(4/7) = (1*4)/(4*7) = 4/28 = 1/7

jimthompson5910 (jim_thompson5910):

so the final answer is 1/7

OpenStudy (anonymous):

Omg with my mind doing sifferent stuff it's so hard to concentrate... Idk how i'm gonna pass..

jimthompson5910 (jim_thompson5910):

go through this whole thing again and let me know if you have any questions

jimthompson5910 (jim_thompson5910):

you saw how I got 1/4 for the first part right?

OpenStudy (anonymous):

Ugh x-x no :(

jimthompson5910 (jim_thompson5910):

ok just say so and I'll go over it again

jimthompson5910 (jim_thompson5910):

there are 8 bills total there are 2 bills that are tens so the probability of picking a ten is 2/8 = 1/4 with me so far?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

once you pick a bill on the first selection, there are 8 - 1 = 7 bills left over there are 4 one dollar bills, out of 7 total (leftover), so the probability of selecting a one dollar bill on the next draw is 4/7 still making sense?

OpenStudy (anonymous):

Yush

jimthompson5910 (jim_thompson5910):

ok good

jimthompson5910 (jim_thompson5910):

you then multiply the two fractions: 1/4 and 4/7

jimthompson5910 (jim_thompson5910):

(1/4)*(4/7) = (1*4)/(4*7) (1/4)*(4/7) = 4/28 (1/4)*(4/7) = 1/7 still with me?

OpenStudy (anonymous):

yuss

jimthompson5910 (jim_thompson5910):

ok so that's the final answer because you just multiply the two probabilities

OpenStudy (anonymous):

so the answer is 1/7?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

umm okay...

jimthompson5910 (jim_thompson5910):

where are you stuck

OpenStudy (anonymous):

No where it's fine

jimthompson5910 (jim_thompson5910):

alright